5 m wire made of gold ( ρgold = 2.4 x 10−8 ) has a cross-sectional area of 2 mm.





Calculate the voltage if a 32 A current runs through the wire.

1
Expert's answer
2022-06-17T08:14:53-0400

The Ohm's law says

"I=\frac{V}{R}"

Hence, the voltage

"V=IR=I\rho\frac{l}{A}"

"V=32*2.4*10^{-8}*\frac{5}{3.14*0.001^2}=1.2\:\rm V"